\begin{thebibliography}{AKMW02}

\bibitem[AKMW02]{MR2003c:14016}
Dan Abramovich, Kalle Karu, Kenji Matsuki, and Jaros{\l}aw W{\l}odarczyk.
\newblock Torification and factorization of birational maps.
\newblock {\em J. Amer. Math. Soc.}, 15(3):531--572, 2002.

\bibitem[Bat99]{MR2000i:14059}
Victor~V. Batyrev.
\newblock Birational {C}alabi-{Y}au {$n$}-folds have equal {B}etti numbers.
\newblock In {\em New trends in algebraic geometry (Warwick, 1996)}, volume 264
  of {\em London Math. Soc. Lecture Note Ser.}, pages 1--11. Cambridge Univ.
  Press, Cambridge, 1999.

\bibitem[BL]{math.AG/0206241}
Lev Borisov and Anatoly Libgober.
\newblock {McKay correspondence for elliptic genera}, arXiv:math.AG/0206241.
\newblock {T}o appear in the Annals of Math.

\bibitem[BL03]{MR1953295}
Lev Borisov and Anatoly Libgober.
\newblock Elliptic genera of singular varieties.
\newblock {\em Duke Math. J.}, 116(2):319--351, 2003.

\bibitem[BS81]{MR83h:32011}
J.-P. Brasselet and M.-H. Schwartz.
\newblock Sur les classes de {C}hern d'un ensemble analytique complexe.
\newblock In {\em The Euler-Poincar\'e characteristic (French)}, volume~83 of
  {\em Ast\'erisque}, pages 93--147. Soc. Math. France, Paris, 1981.

\bibitem[DL92]{MR93g:11118}
J.~Denef and F.~Loeser.
\newblock Caract\'eristiques d'{E}uler-{P}oincar\'e, fonctions z\^eta locales
  et modifications analytiques.
\newblock {\em J. Amer. Math. Soc.}, 5(4):705--720, 1992.

\bibitem[DL99]{MR99k:14002}
Jan Denef and Fran{\c{c}}ois Loeser.
\newblock Germs of arcs on singular algebraic varieties and motivic
  integration.
\newblock {\em Invent. Math.}, 135(1):201--232, 1999.

\bibitem[DL01]{MR1905328}
Jan Denef and Fran{\c{c}}ois Loeser.
\newblock Geometry on arc spaces of algebraic varieties.
\newblock In {\em European Congress of Mathematics, Vol. I (Barcelona, 2000)},
  volume 201 of {\em Progr. Math.}, pages 327--348. Birkh\"auser, Basel, 2001.

\bibitem[Ful84]{MR85k:14004}
William Fulton.
\newblock {\em Intersection theory}.
\newblock Springer-Verlag, Berlin, 1984.

\bibitem[Ken90]{MR91h:14010}
Gary Kennedy.
\newblock Mac{P}herson's {C}hern classes of singular algebraic varieties.
\newblock {\em Comm. Algebra}, 18(9):2821--2839, 1990.

\bibitem[Kol92]{MR94f:14013}
J.~Koll\'ar.
\newblock {\em Flips and abundance for algebraic threefolds}.
\newblock Soci\'et\'e Math\'ematique de France, Paris, 1992.
\newblock Papers from the Second Summer Seminar on Algebraic Geometry held at
  the University of Utah, Salt Lake City, Utah, August 1991, Ast\'erisque No.
  211 (1992).

\bibitem[Lo02]{MR2003k:14010}
Eduard Looijenga.
\newblock Motivic measures.
\newblock {\em Ast\'erisque}, (276):267--297, 2002.
\newblock S\'eminaire Bourbaki, Vol.\ 1999/2000.

\bibitem[LW90]{MR91g:32039}
Anatoly~S. Libgober and John~W. Wood.
\newblock Uniqueness of the complex structure on {K}\"ahler manifolds of
  certain homotopy types.
\newblock {\em J. Differential Geom.}, 32(1):139--154, 1990.

\bibitem[Mac74]{MR50:13587}
Robert~D. MacPherson.
\newblock Chern classes for singular algebraic varieties.
\newblock {\em Ann. of Math. (2)}, 100:423--432, 1974.

\bibitem[Wan03]{MR2003j:14015}
Chin-Lung Wang.
\newblock {$K$}-equivalence in birational geometry and characterizations of
  complex elliptic genera.
\newblock {\em J. Algebraic Geom.}, 12(2):285--306, 2003.

\bibitem[Vey03]{MR2030094}
Willem Veys.
\newblock Stringy zeta functions for {$\mathbb Q$}-{G}orenstein varieties.
\newblock {\em Duke Math. J.}, 120(3):469--514, 2003.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsfonts,latexsym,amsthm,amssymb}
\usepackage[all]{xy}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corol}[theorem]{Corollary}
\newtheorem{prop}[theorem]{Proposition}

{\theoremstyle{definition} }
{\theoremstyle{remark} 
}

\newcommand{\Abb}{{\mathbb{A}}}
\newcommand{\Cbb}{{\mathbb{C}}}

\newcommand{\Lbb}{{\mathbb{L}}}
\newcommand{\Pbb}{{\mathbb{P}}}
\newcommand{\Qbb}{{\mathbb{Q}}}
\newcommand{\Zbb}{{\mathbb{Z}}}

\newcommand{\Til}[1]{{\widetilde{#1}}}

\DeclareMathOperator{\Spec}{Spec}

\newcommand{\one}{1\hskip-3.5pt1}
\newcommand{\csm}{c_{\text{\rm SM}}}

\title{Chern classes of birational varieties}
\author{Paolo Aluffi}
\address{Florida State University, Tallahassee, Florida}
\email{aluffi@math.fsu.edu}

\begin{document}

\begin{abstract}
Let $\varphi: V\dashrightarrow W$ be a birational map between smooth
algebraic varieties which does not change the canonical class (in the
sense of Batyrev, \cite{MR2000i:14059}). We prove that the total
homology Chern classes of $V$ and $W$ are push-forwards of the same
class from a resolution of indeterminacies of $\varphi$.

For example, it follows that the push-forward of the total Chern class
of a crepant resolution of a singular variety is independent of the
resolution.
\end{abstract}

\maketitle

\section{Introduction and statement of the result}\label{intro}

There is a strong motivic feel to the theory of
Chern-Schwartz-MacPherson classes, although this does not seem to have
yet congealed into a precise statement in the literature. In this note
we give an instance of a result exploiting this heuristic observation.
Our statement is an analog of Victor Batyrev's well-known theorem
showing that the Betti numbers of birational manifolds `in the same
$K$-class' are equal, see \cite{MR2000i:14059}. Batyrev's theorem has
been generalized to many other numerical invariants, such as
the Hodge numbers or certain Chern numbers. Our statement is not
numerical; it refers to the total Chern class of the tangent bundle, a
class in the Chow group of the variety.

\begin{theorem}\label{main}
Let $\varphi: V \dashrightarrow W$ be a birational morphism of
nonsingular algebraic varieties over an algebraically closed field of
characteristic~0. Assume that there is a resolution of
indeterminacies of $\varphi$, $Z$:
$$\xymatrix{
& Z \ar[ld]_v \ar[rd]^w \\
V \ar@{-->}[rr] & & W
}$$
such that $v$, $w$ are proper and birational, and the Jacobian ideals
of $v$ and $w$ coincide. 

Then there exists a class $C\in (A_*Z)_\Qbb$ such that
$$c(TV)\cap [V] = v_*(C)\quad\text{and}\quad c(TW)\cap [W]= w_*(C)$$
in $(A_*V)_\Qbb$, $(A_*W)_\Qbb$ respectively.
\end{theorem}

The condition on the Jacobians implies that the pull-backs $v^* K_V$
and $w^* K_W$ of the respective canonical classes coincide. In fact,
as Lawrence Ein kindly pointed out to me, the condition {\em is
equivalent\/} to requiring that $v^* K_V \equiv_{\text{Num}} w^* K_W$,
an easy consequence of the `Kodaira lemma' (see for example 2.19 in
\cite{MR94f:14013}). 

Theorem~\ref{main} has a number of immediate consequences: for instance
the equality of Euler characteristics of $V$ and $W$, or more generally
the equality of Chern numbers
$$c_1(V)^i\cdot c_{n-i}(V)=c_1(W)^i\cdot c_{n-i}(W)$$
for all $i\ge 0$ (where $n=\dim V=\dim W$) in the stated hypotheses,
when $V$ and $W$ are complete. Identities of this type are not new;
for example, Anatoly Libgober and John Wood have proved
(\cite{MR91g:32039}) that $c_1\cdot c_{n-1}$ is determined by the
Hodge numbers, and these agree for $V$ and $W$, as mentioned above. 
Our argument is very direct, and yields the equality for all $i$;
suprisingly, this simple equality does not seem to be explicitly in
the literature for $i\ge 2$, although it is certainly implied by 
recent and powerful results on elliptic genera (\cite{MR1953295},
\cite{MR2003j:14015}).  Also:

\begin{corol}\label{crep}
Let $\alpha: Y \to X$ be a crepant resolution. Then the class
$$\alpha_*(c(TY)\cap [Y])$$
in $(A_*X)_\Qbb$ is independent of $Y$.
\end{corol}

Again, this is immediate from Theorem~\ref{main}; it also follows from
the more sophisticated technology in Lev Borisov and Libgober's work. 
It would be interesting to extend this notion of `Chern class' to
arbitrary singular varieties, and to compare it to other known
intrinsic definitions such as the Chern-Fulton class
(\cite{MR85k:14004}, Example~4.2.6), or the Chern-Schwartz-MacPherson
class (\cite{MR50:13587}).

Similarly to analogous numerical statements, Theorem~\ref{main} is an
easy consequence of a basic formula in motivic integration, to wit
Proposition~6.3.2 in \cite{MR99k:14002}. This formula is usually
obtained as a corollary of Kontsevich's change of variable formula; we
refer the reader to loc.~cit.~or to the excellent surveys
\cite{MR1905328}, \cite{MR2003k:14010} for a discussion in the context
of motivic integration. 

In order to provide a self-contained treatment of the material, we
take the liberty of providing an alternative proof of this `motivic'
statement, using the factorization theorem of \cite{MR2003c:14016};
our argument has the (very small) advantage of proving the needed
formula without requiring completions. This is done in
\S\ref{motiv}. The idea of using the factorization theorem as an
alternative to motivic integration is of course not new; for example,
Willem Veys uses it to prove a similar (in fact, stronger) statement,
in \S2 of \cite{MR2030094}. Also, the factorization theorem is a
key tool in the work of Borisov and Libgober on elliptic genera. In
fact, Borisov has pointed out to me that Theorem~\ref{main} can also
be derived from Theorem~3.5 in \cite{math.AG/0206241}, and that this
work may be used to extend the class given in Corollary~\ref{crep} to
all varieties with log-terminal singularities.

Regardless of how the basic formula is established, however, our main
observation is that this formula alone---and in fact even just the
Euler characteristic version of the basic formula, which goes back to
\cite{MR93g:11118}---implies an analog at the level of
Chern-Schwartz-MacPherson classes, and that this immediately implies
Theorem~\ref{main}. The upgrade of the basic formula to
Chern-Schwartz-MacPherson classes is Theorem~\ref{strmain};
Theorem~\ref{main} follows easily from this result, as shown in
\S\ref{CSM}.

Theorem~\ref{strmain} must thus have been known to the experts very
early (and Fran\c cois Loeser confirms this); the fact that it has
nice applications, such as the invariance of $c_1^i\cdot c_{n-i}$
under $K$-equivalence, seems to have been inexplicably overlooked thus
far. It may be argued that Theorem~\ref{main}, and hence these
straightforward consequences, should have been immediately noticed
after the diffusion of \cite{MR93g:11118}; in particular, long before
the discovery of the factorization theorem of \cite{MR2003c:14016},
which has fueled the more recent and sophisticated work in this
direction. 

In any case, we believe that Theorem~\ref{strmain} is of independent
interest in the context of Chern-Schwartz-MacPherson classes, and that
it deserves to be out in the open.

I thank Prof.~Kenji Matsuki for promptly answering many queries
concerning \cite{MR2003c:14016}. 
Thanks are also due to the Max-Planck-Institut f\"ur Mathematik in 
Bonn, where part of this work was done.

\section{The basic formula}\label{motiv}

We work over an algebraically closed field $k$. We will assume $k$ has
characteristic~0, as this is at present needed for resolution of
singularities (needed in the factorization theorem of
\cite{MR2003c:14016}) and the theory of Chern-Schwartz-MacPherson
classes used in \S\ref{CSM}.

If $X$ is a variety, $[X]$ will denote its class in the Grothendieck
ring of algebraic varieties over $k$; that is, the abelian group
generated by isomorphism classes of varieties modulo the relation
$[X-Y]=[X]-[Y]$ whenever $Y$ is a closed subvariety of $X$, with
multiplication defined by product over~$k$. The unit element is
$1=[\Spec k]$. It is costumary to denote by $\Lbb$ the `Tate motive'
$[\Abb^1]$.

The basic formula is a relation in this ring, depending on the datum
of a `modification' of nonsingular varieties. It is in fact
convenient to localize the ring at elements
$[\Pbb^\mu]=\frac{\Lbb^{\mu+1}-1}{\Lbb-1}$, and this will be done
without further mention in what follows.

{\bf Notation:} if $\{D_j\}_{j\in J}$ is a set of irreducible divisors
in a variety and $I\subset J$, then $D_I^\circ$ will denote the
complement of $\cup_{j\in J\setminus I}D_j$ in $\cap_{i\in I}D_i$.

Let $v: Z \to V$ be a proper birational morphism of nonsingular
varieties, such that the exceptional divisor of $v$ has normal
crossings, with nonsingular irreducible components $E_j$, $j\in J$.

Assume that the Jacobian ideal of $v$ is principal, with divisor
$\sum_j \mu_j E_j$, so that $v^*K_V=K_Z+\sum_j \mu_j E_j$. The
following formula is (essentially) a particular case of Proposition~6.3.2 
in \cite{MR99k:14002}.

\begin{theorem}\label{basic}
Let $U$ be a closed subvariety of $V$. Then
$$[U]=\sum_{I\subset J}\frac{[E_I^\circ\cap v^{-1}(U)]}{\prod_{i\in I}
[\Pbb^{\mu_i}]}\quad.$$
\end{theorem}

As mentioned in \S\ref{intro}, we provide a motivic-integration-free
proof of Theorem~\ref{basic} in this section, which the hurried reader
may safely skip.

The factorization theorem of \cite{MR2003c:14016} will reduce the
proof of Theorem~\ref{basic} to the case in which $v$ is a blow-up at
a smooth center; this case will be dealt with by analyzing the
situation in the exceptional divisor, and this in turn will be reduced
to the analysis of a single fiber, a projective space. The real reason
behind Theorem~\ref{basic} can then be traced to the following simple
lemma, whose statement requires no localization.

\begin{lemma}\label{simplex}
Let $d\ge k>0$ and $\mu_i\ge 0$, $i=1,\dots, k$, be integers, and let
$K=\{1,\dots,k\}$. Let $H_1,\dots,H_k$ be linearly independent
hyperplanes in $\Pbb^{d-1}$. 
Then
$$\sum_{I\subset K} \left([H_I^\circ]\prod_{i\in K\setminus 
I}[\Pbb^{\mu_i}]\right)=[\Pbb^{\sum_{j\in K}\mu_j+d-1}]\quad.$$
\end{lemma}

\begin{proof}
With evident coordinates, the subset $H_I^\circ$ consists of
homogeneous $d$-tuples in which $|I|$ fixed components are~0 and
$k-|I|$ fixed components are nonzero. It follows that
$$[H_I^\circ]=\left\{
\aligned
\frac{(\Lbb-1)^{k-|I|}\Lbb^{d-k}}{\Lbb-1} &\quad |I|\ne k\\
\frac{\Lbb^{d-k}-1}{\Lbb-1}  \qquad &\quad |I|=k
\endaligned
\right.\quad,$$
giving
$$[H_I^\circ]\prod_{i\in K\setminus I}[\Pbb^{\mu_i}]=\left\{
\aligned
\frac{\Lbb^{d-k}}{\Lbb-1} \prod_{i\in K\setminus I}(\Lbb^{\mu_i+1}-1) 
&\quad I\ne K\\
\frac{\Lbb^{d-k}-1}{\Lbb-1} - \frac1{\Lbb-1} \quad &\quad I=K
\endaligned
\right.\quad.$$
Hence the left-hand-side in the stated equality is
$$\frac{\Lbb^{d-k}}{\Lbb-1}\sum_{I\subset K} \prod_{i\in K\setminus I} 
(\Lbb^{(\mu_i+1)}-1) - \frac 1{\Lbb-1}
= \frac{\Lbb^{d-k}}{\Lbb-1}\prod_{j\in K} (1+ (\Lbb^{\mu_j+1}-1)) - 
\frac 1{\Lbb-1}$$
which immediately yields the right-hand-side.
\end{proof}

The form in which Lemma~\ref{simplex} will be used is the following.
\begin{corol}\label{simplexcor}
With the same notation as in Lemma~\ref{simplex}, set 
$\mu_0=\sum_{j\in K}\mu_j+d-1$; then
$$\sum_{I\subset K} \frac{[H_I^\circ]}
{\prod_{i\in (\{0\}\cup I)}[\Pbb^{\mu_i}]}
=\frac 1{\prod_{j\in K}[\Pbb^{\mu_j}]}\quad.$$
\end{corol} 

This formula will be used to control the situation across blow-ups. 
Let $X$ be a nonsingular algebraic variety, and assume given a set
$\{E_j\}_{j\in J}$ of irreducible nonsingular divisors in $X$. Let
$S\subset X$ be a nonsingular subvariety of codimension~$d\ge 1$,
intersecting $\cup E_j$ with normal crossings: that is, at each point
$s\in S$ there is an analytic system of parameters $x_1,\dots,x_n$ for
$X$ at $s$ such that $S$ is given by $x_1=\dots=x_d=0$, and $\cup E_j$
is given by a monomial in the $x_i$'s. 

Let $\pi:Y\to X$ be the blow-up of $X$ along $S$; denote by $\Til E_0$
the exceptional divisor $\pi^{-1}(S)$, and let $\Til E_j$ be the
proper transform of $E_j$. The following lemma is a straightforward
computation, which we leave to the reader.

\begin{lemma}\label{standard}
---The divisor $\Til E_0\cup(\cup_{j\in J} \Til E_j)$ has normal
crossings at all points of $\Til E_0$;

---If $E_j$ does not contain $S$ and $s\in S\cap E_j$, then $\Til
E_j$ contains the fiber $\pi^{-1}(s)$;

---Let $K$ denotes the set of indices $j\in J$ such that $E_j$
contains $S$. Then for all $s\in S$ and $j\in K$, the intersections of
$\Til E_j$ with $\pi^{-1}(s)\cong \Pbb^{d-1}$ consist of linearly
independent hyperplanes in $\Pbb^{d-1}$.
\end{lemma}

Now assign to each $E_j$ and $\Til E_j$, $j\in J$ a nonnegative integer
$\mu_j$, and assign to the exceptional divisor $\Til E_0$ the combination
$$\mu_0:=\sum_{E_j\supset Z}\mu_j+d-1\quad.$$
Note that, with this choice,
$$(\sum_{j\in (\{0\}\cup J)}\mu_j \Til E_j) - \pi^* (\sum_{j\in J}
\mu_j E_j)$$
is the divisor of the Jacobian ideal of $\pi$.

For any algebraic subset $T\subset X$, let
$$\chi_X(T):=\sum_{I\subset J}\frac{[E_I^\circ\cap T]}
{\prod_{i\in I}[\Pbb^{\mu_i}]}$$
and
$$\chi_Y(\pi^{-1}(T)):=\sum_{I\subset (\{0\}\cup J)}\frac
{[\Til E_I^\circ\cap \pi^{-1}(T)]}{\prod_{i\in I}[\Pbb^{\mu_i}]}
\quad.$$

\begin{prop}\label{pre}
With notation as above, $\chi_Y(\pi^{-1}(T))=\chi_X(T)$.
\end{prop}

\begin{proof}
The contributions to $\chi_X(T)$ coming from the complement of $S$
equal the contributions to $\chi_Y(\pi^{-1}(T))$ coming from the
complement of $\Til E_0$. Thus it suffices to prove that 
$$\sum_{I\subset J}\frac{[E_I^\circ\cap T\cap S]}{\prod_{i\in I}
[\Pbb^\mu_i]}$$ 
equals the contribution to $\chi_Y(\pi^{-1}(T))$ supported
within $\Til E_0$, which consists of
$$\sum_{I\subset J}\frac{[\Til E_{\{0\}\cup I}^\circ\cap
  \pi^{-1}(T)]}{[\Pbb^{\mu_0}]\prod_{i\in I}[\Pbb^\mu_i]}\quad.$$
Now every point of $T\cap S$ belongs to exactly one $E_K^\circ$.
For all $K\subset J$, $\Til E_{\{0\}\cup K}^\circ\cap
  \pi^{-1}(T)$ fibers onto $E_K^\circ\cap T\cap S$ with fibers equal
to the union over all $I\subset K$ of
$$\Til E_I^\circ\cap \pi^{-1}(x) \cong \Til E_I^\circ\cap \Pbb^{d-1}$$
for $x\in E_K^\circ\cap T\cap S$ (identifying the fiber of $\Til E_0$
with $\Pbb^{d-1}$). By the multiplicativity of products in the
Grothendieck ring, it suffices to show that for all $K\subset J$
$$\sum_{I\subset K}\frac{[\Til E_I^\circ \cap \pi^{-1}(x)]}
{[\Pbb^{\mu_0}]\prod_{i\in I}[\Pbb^{\mu_i}]} = \frac 1{\prod_{j\in K}
  [\Pbb^{\mu_j}]}\quad,$$
for all $x\in T\cap S\cap E_K^\circ$. By Lemma~\ref{standard}, indices
in $K$ corresponding to divisors which do not contain $S$ only
contribute common multiplicative factors to the denominators in both
sides of this equality; so we may assume that, for all $j$ in $K$,
$S\subset E_j$.

Again by Lemma~\ref{standard}, the intersections $\pi^{-1}(x)\cap \Til
E_j$ consist of linearly independent hyperplanes in $\pi^{-1}(x)\cong
\Pbb^{d-1}$; with the given $\mu_0$, the sought formula is then
precisely the statement of Corollary~\ref{simplexcor}.
\end{proof}

We are now ready to prove Theorem~\ref{basic}. 

\begin{proof}
By the factorization theorem of \cite{MR2003c:14016}, $v$ may be
decomposed as a sequence:
$$\xymatrix{
Z=V_0 \ar@{-->}[r] & V_1 \ar@{-->}[r] & V_2 \ar@{-->}[r] & \dots
\ar@{-->}[r] & V_M=V
}$$
where each map is either a blow-up or a blow-down at a smooth center.
More precisely, for each $i$ one has one of the possibilities
$$\xymatrix{
V_i \ar[rr] \ar[rd] & & V_{i+1} \ar[ld] \\
& V
}\qquad
\xymatrix{
V_{i+1} \ar[rr] \ar[rd] & & V_i \ar[ld] \\
& V
}$$
where the diagonal (regular!) maps are obtained by composing the 
rational maps to $V$; and the horizontal map is a blow-up of a
nonsingular variety at a nonsingular center. This center may be chosen
so that it meets with normal crossings the exceptional divisor of the
map to $V$ (cf.~\cite{MR2003c:14016}, expecially the `Furthermore'
section of the main statement, and part~6 of Theorem~0.3.1).

Therefore, at each stage of the factorization we are in the hypotheses
of Proposition~\ref{pre}, with the normal crossing divisor on the base
of the blow-up equal to the Jacobian of the map to $V$; this is
compatible with the prescribed multiplicities $\mu_j$, since the
composition of the differentials is the differential of the
composition. By pasting together the equalities thereby obtained at
each stage, we get that $\chi_Z(\pi^{-1}(U))=\chi_V(U)=[U]$; that is,
the stated equality.
\end{proof}

\section{Proof of the theorem}\label{CSM}

A {\em constructible function\/} on a variety $Z$ is a finite
$\Zbb$-linear combination of characteristic functions of
subvarieties: $\sum_{S\subset Z} m_S \one_S$ with $m_S\in \Zbb$, and
$\one_S(s)=1$ for $s\in S$, $\one_S(s)=0$ otherwise. 

Denote by $F(Z)$ the group of constructible functions on the
variety~$Z$. This assignment is in fact a covariant functor under
proper morphisms, as follows: if $v: Z \to V$ is a proper morphism,
and $S$ is a subvariety (or more generally a constructible set) in
$Z$, define the push-forward of the function $\one_S$ by 
$$\forall p\in V\quad,\quad v_*(\one_S)(p)=\chi(v^{-1}(p)\cap
S)\quad.$$
Here $\chi$ denotes the topological Euler characteristic if the ground
field is $\Cbb$; this definition may be extended to arbitrary
algebraically closed fields of characteristic~0, see
\cite{MR91h:14010}.

In \cite{MR50:13587}, Robert MacPherson proved (over $\Cbb$) that
there exists a natural transformation $c$ from $F$ to homology, such
that if $V$ is a {\em nonsingular\/} variety, then $c(\one_V)$ equals
the total homology Chern class of the tangent bundle of $V$,
$c(TV)\cap [V]$. MacPherson's natural transformation can in fact be
lifted to the Chow group and over any algebraically closed field of
characteristic~0, see Example~19.1.7 in \cite{MR85k:14004} and
\cite{MR91h:14010}.

MacPherson's natural transformation may be used to define a notion
extending to possibly singular varieties the total Chern class of the
tangent bundle. This notion is commonly named
`Chern-Schwartz-MacPherson class', since it can be shown that the
class so obtained agrees with a notion previously defined by
Marie-H\'el\`ene Schwartz, cf.~\cite{MR83h:32011}.

Analogously, we may define a `Chern class' of any constructible set
$S$ in a variety $Z$, {\em as an element of $A_*Z$,\/} by applying
MacPherson's transformation to its characteristic function: that is,
we set 
$$\csm(S):=c(\one_S)\in A_*Z\quad.$$

Now consider the situation of \S\ref{motiv}: $v:Z\to V$ is a proper
birational morphism of nonsingular varieties; the exceptional divisor
has normal crossings, with nonsingular components $E_j$, and the
Jacobian ideal of $v$ is principal, with divisor $\sum_j \mu_j E_j$.

\begin{theorem}\label{strmain}
Let $U$ be a constructible subset of $V$. Then
$$\csm(U)=v_* \sum_{I\subset J}\frac{\csm(E_I^\circ\cap
  v^{-1}(U))}{\prod_{i\in I}(\mu_i+1)}$$
in $(A_*V)_\Qbb$.
\end{theorem}

\begin{proof}
For all $p\in V$,
$$\sum_{I\subset J} \frac{[E_I^\circ\cap v^{-1}(p)]}{\prod_{i\in I}
  [\Pbb^{\mu_i}]}=[p]$$
by Theorem~\ref{basic}. Taking Euler characteristics:
$$\sum_{I\subset J} \frac{\chi(E_I^\circ\cap v^{-1}(p))}{\prod_{i\in I}
  (\mu_i+1)}=1\quad,$$
thus,
$$\sum_{I\subset J} \frac{\chi(E_I^\circ\cap v^{-1}(U)\cap
  v^{-1}(p))}{\prod_{i\in I} (\mu_i+1)}=
\left\{\aligned
0 & \quad p\not\in U\\
1 & \quad p\in U
\endaligned\right.
\quad.$$
By definition of push-forward of constructible functions, this tells
  us that
$$v_*\sum_{I\subset J} \frac{\one_{E_I^\circ\cap v^{-1}(U)}}
{\prod_{i\in I}(\mu_i+1)} =\one_U$$
in $F(V)\otimes\Qbb$. Applying MacPherson's natural transformation
gives the statement.
\end{proof}

Theorem~\ref{main} follows immediately. Indeed, if $V$ and $W$ are
both dominated birationally by~$Z$, and $v: Z \to V$, $w: Z \to W$
have the same Jacobian, we may assume (after resolution of
singularities) that $Z$ is nonsingular, with normal crossing
exceptional divisor $\cup E_j$, and both $v$ and $w$ have Jacobian
with divisor $\sum \mu_j E_j$. Let
$$C=\sum_{I\subset J}\frac{\csm(E_I^\circ)}{\prod_{i\in I}(\mu_i+1)}
\in (A_*Z)_\Qbb\quad;$$
then $\csm(V)=v_*(C)$, $\csm(W)=w_*(C)$ by Theorem~\ref{strmain}, and
these classes agree with $c(TV)\cap [V]$, $c(TW)\cap [W]$
respectively, by the normalization property of
Chern-Schwartz-MacPherson's classes.

Theorem~\ref{strmain} appears to be of independent interest. For
example, let $U$ be a singular variety embedded in a nonsingular
variety $V$, and let $Z$ be the variety obtained from $V$ as $U$
undergoes a resolution of singularities \`a la Hironaka. Then
Theorem~\ref{strmain} gives an explicit expression for the
Chern-Schwartz-MacPherson class of $U$, with rational coefficients, 
in terms of classes of loci in its inverse image in $Z$ (note that
in this case $E_I^\circ\subset v^{-1}(U)$ for $I\ne\emptyset$). 
To our knowledge, this expression is new.

\end{document}